# Strong Decays of the Light Exotic $0^{+-}$ and $2^{+-}$ Hybrid Mesons

### v2

@@ Abstract
A model of hybrid meson structure based on the QCD Hamiltonian in Coulomb gauge and the use of a single constituent quasigluon is applied to the computation of hadronic decays of mesons with exotic quantum numbers, $0^{+-}$ and $2^{+-}$. These correspond to hybrid mesons in which the gluon couples to a $q\bar{q}$ pair in an $P$-wave and can therefore be identified as orbital excitations of the exotic $1^{-+}$ state. Interestingly, we find $0^{+-}$ states to be narrow, contrary to what was found by previous calculations. This is primarily due to the suppression of the decay mode $0^{+-}\rightarrow 2^{1}S_{0}+1^{1}S_{0}$, which is unique to our model. The $2^{+-}$ states are found to be narrow, as expected.
@@ I Introduction
Indeed, several experimental efforts such as the GlueX at JLab, PANDA at FAIR and BESIII at BEPC II will be, or are currently under way, and there is hope that future data, combined with continuous progress in lattice field computations, will shed light on these enigmatic states. However, the field is not without progress. The puzzle of the two exotic $1^{-+}$ states labeled $\pi_{1}(1400)$ and $\pi_{1}(1600)$ Workman:2022ynf seems to have mainly been resolved, suggesting that only the heavier of the two is necessary to explain the data JPAC:2018zyd ; Kopf:2020yoa . Just recently, the BESIII collaboration BESIII:2022riz has observed an isoscalar resonance with exotic quantum numbers $1^{-+}$ in $J/\psi\to\gamma\eta\eta^{\prime}$. This new state, called the $\eta_{1}(1855)$, has a mass of $1855\pm 9{}^{+6}_{-1}$ MeV and width of $188\pm 18{}^{+3}_{-8}$ MeV, making it the ideal candidate for the isoscalar partner to the isovector $\pi_{1}(1600)$ BESIII:2022riz ; Swanson:2023zlm .
@@
Models of hybrid mesons have a long history, reaching back to the 1970s, and they can be broadly divided into two categories: flux-tube models and constituent gluon models (see Meyer:2015eta for a review). These two camps primarily differ in how they describe the gluonic degrees of freedom. In flux-tube models flux (which include string models string ), glue is represented as a collective string-like excitation, while in constituent models the gluonic degrees of freedom are described by quasiparticles. The latter notion finds confimation in recent lattice studies demonstrating that the low-lying charmonium hybrid spectrum is well-described by a single gluonic quasiparticle with axial quantum numbers $J_{G}^{PC}=1^{+-}$ Dudek:2011bn .
@@
Throughout the years, constituent gluon models have developed from almost dynamics-free models HM into more modern pictures which include many of the features of QCD Swanson:1997wy ; Swanson:1998kx ; Guo:2008yz . More recently, the constituent picture with the QCD Hamiltonian in Coulomb gauge as a starting point has been applied to charmonium hybrids decays Farina:2020slb , the light $H_{1}$ states ($1^{--}$, $(0,1,2)^{-+}$) Farina:2023oqk , and to a study of light hybrid meson mixing Swanson:2023zlm . In this paper, we use this picture to compute the strong decays of the light exotic $0^{+-}$, $2^{+-}$ and $2^{+-^{\prime}}$ hybrid mesons. These correspond to hybrid mesons in which the gluon couples to a $q\bar{q}$ pair in an $P$-wave and can therefore be identified as the orbital excitations of the exotic $1^{-+}$ state previously studied in Farina:2023oqk .
@@
The paper is organized as follows: Section II contains details of the QCD Hamiltonian in Coulomb Gauge, which forms the basis of our model, and the construction of the light hybrid meson state from constituent quasiparticles. Model parameters used in the computation and the decay model are also contained in this section, along with the resulting selection rules. Section III presents the results obtained for both isoscalar and isovector hybrids. Summary and conclusions, along with prospects for potential discovery of these states, are found in Section IV.
@@ II.1 QCD Hamiltonian In Coulomb Gauge
The Coulomb kernel is instantaneous and contains all powers of gluonic field operators. These may be organized in a tower of Dyson equations and truncations can be solved. This approach shows that confinement appears with the aid of simple variational vacuum ansatzSzczepaniak:2001rg . However, this potential need not be equivalent to the interaction between static color sources, which is typically used in phenomenological calculationsDawid:2024uqz . We therefore choose to simplify by replacing the vacuum matrix element of the Coulomb interaction with:
@@
where $a_{s}$, $b$, and $\mathcal{C}$ are parameters to be determined from fitting to the meson spectrum, as explained below. This choice, besides reproducing the well-known Cornell potential of non-relativistic meson models, finds support in lattice simulations of the $q\bar{q}$ static quark potentials with the gluonic field in different configurations, which show that they all exhibit a long-range linear behavior (see, for example, Ref. Juge:1997nc ).
@@
where $m_{g}$ is the dynamical gluon mass given by $m_{g}\approx 600$ MeV and $b_{g}$ is the gluon parameter with value $b_{g}\approx 6000$ MeV. We remark that despite the appearance of an effective mass, the quasigluon remains transverse.
@@ II.2 Light Hybrids
The hybrid meson state is constructed, by assumption, from the minimal number of constituent quasiparticles: a $q\bar{q}$ pair coupled with an axial gluon. This assumption is made plausible here by the relatively large gluon mass. The total meson angular momentum $J$ is obtained by first coupling the gluon spin projection to the gluon angular momentum, $\ell_{g}$, to obtain the total gluon angular momentum $j_{g}$. Calculations can be streamlined by working in the gluon helicity basis. Doing so and assuming that $\ell_{g}=j_{g}$ produces a Clebsch-Gordan factor of
@@
This represents a transverse electric (TE) gluon and forms the explicit realization of the axial constituent gluon. Setting $\ell_{g}=j_{g}\pm 1$ yields a transverse magnetic (TM) gluon with a Clebsch factor given by $\chi^{(+)}_{\lambda,\mu}=\delta_{\lambda,\mu}/\sqrt{2}$. Combining with quark spins yields the final generic expression for a hybrid creation operator
@@
Since all lowest lying hybrids have $\xi=-1$ and $j_{g}=1$, only the TE mode is of interest for this work. Furthermore, the hybrid state is an eigenstate of parity and charge conjugation with eigenvalues given by
@@ II.3 Model Parameters
The spectrum can be obtained by expanding the hybrid wavefunction in Gaussian basis, which was previously used to compute the spin-averaged spectrum of charmonium hybrids Farina:2020slb . Here, however, since the main goal is to compute the decay properties of the mesons, we are only interested in the gross features of meson structure and the resulting hybrid wavefunctions are therefore approximated by a product of two Gaussians with widths $\alpha$ and $\beta$:
@@
Similarly, ordinary mesons are described by SHO wavefunctions depending on a scale parameter $\beta_{m}$.
@@ II.4 Decay Model
This decay model yields two selection rules that are typical of hybrid decay models, namely: (i) spin zero hybrids do not decay to spin zero mesons, (ii) TE hybrids do not decay to mesons with identical spatial wavefunctions. The former follows from the $^{3}S_{1}$ quantum numbers of the vertex, while the latter follows from the symmetry properties of the decay amplitude (see Farina:2020slb for details). Additionally, decays to an $S$-wave and $P$-wave meson pair are generally dominant.
@@ III.1 Hybrid Masses
Hybrid masses constitute another source of uncertainty in the model. Because our model only gives spin-averaged hybrid masses, and there are currently no experimental data to constrain them, one must rely on lattice data. However, the results from lattice gauge theory also have variability. For example, a study of the lightest hybrids supermultiplet Dudek:2011bn computes isovector masses of all the $H_{1}$ hybrids and the exotic $0^{+-}$ and $2^{+-}$ hybrids at different pion masses, ranging from $m_{\pi}\approx 700$ MeV down to $m_{\pi}\approx 400$ MeV. A crude extrapolation down to the physical pion mass gives estimates for the isovector hybrids but not the isoscalars. Alternatively, a more comprehensive study of the full isovector-isoscalar meson spectrum Dudek:2013yja extracted hybrid masses at $m_{\pi}\approx 400$ MeV. The results suggest that there should be a $1^{-+}$ isovector at about $2025$ MeV, and two isoscalar counterparts at higher masses of about $2150$ MeV and $2340$ MeV, respectively. The excited states masses are obtained in the same way, and they show an average mass splitting of about 35 MeV between the isovector and low isoscalar and about 100 MeV between the low and high isoscalar. However, the light isovector $1^{-+}$ mass of about 2025 MeV, is roughly $400$ MeV higher than the accepted $\pi_{1}(1600)$ mass. The approach adopted here, admittedly not overly rigorous, is to shift all of the masses in Ref. Dudek:2013yja down by the same amount (425 MeV) so that the isovector $1^{-+}$ mass matches the $\pi_{1}(1600)$ mass. This keeps the mass splittings the same and gives the masses for the excited states reported in the Table 3.
@@ III.2 Isovectors
$|K_{1L}\rangle$ $=$ $\sin{\theta_{K}}|^{1}P_{1}\rangle+\cos{\theta_{K}}|^{3}P_{1}\rangle$
$|K_{1H}\rangle$ $=$ $-\cos{\theta_{K}}|^{1}P_{1}\rangle+\sin{\theta_{K}}|^{3}P_{1}\rangle.$ (15)
@@
$\Gamma_{tot}$ 22.8 65.9
The results show that the three isovectors are all narrow. While the $2^{+-}$ states were previously reported to be narrow, the $0^{+-}$ comes as a surprise. If we compare our results to those of previous analyses such as the PPS model Page:1998gz (and the IKP model reported therein) or the work of Ref. Chen2025 which uses a very similar approach to ours, the $0^{+-}$ was always reported as being wide, with a total width in the 250-650 MeV range and a dominant contribution of about 200-250 MeV coming from the $\pi(1300)\pi$, which in our work is almost entirely suppressed. In our model, the suppression of the channel is strongly dependent on the value of the $\beta_{m}$ parameter used for the $\pi(1300)$. This dependence is shown in Figure 2, where the total width is plotted as a function of $\beta_{\pi(1300)}$. Oscillatory behavior like the one in the figure are typical of decay channels that include excited states. Remarkably, the preferred value of $\beta_{\pi(1300)}$ lies very close to a node in the decay amplitude, which occurs at about $\beta_{\pi(1300)}=0.195$ GeV. The figure also shows that the decay width rapidly increases with increasing $\beta$, reaching its maximum of about $41.8$ MeV at $\beta_{\pi(1300)}=0.38$ GeV. However, even if we were to use $\beta_{\pi(1300)}=0.38$ GeV, our value for the $\pi\pi(1300)$ width would still be smaller than those of Page:1998gz ; Chen2025 by a factor of at least six.
Additionally, for all three states, $S+P$ wave decays are preferred. This is the case for the $0^{+-}$, whose main decay channels are the $h_{1}(1170)\pi$ and $b_{1}(1235)\eta$, in that order. The same is also true for the $2^{+-}$ and $2^{+-^{\prime}}$ states. However, these states show some differences. The $2^{+-}$ decays mainly into $a_{1}(1260)\pi$, $a_{2}(1320)\pi$ and $h_{1}(1170)\pi$, in that order, but has a small $b_{1}(1235)\eta$ contribution. In contrast, the main decay channels for the $2^{+-^{\prime}}$ are $h_{1}(1170)\pi$ and $a_{1}(1260)\pi$. In this case, $b_{1}(1235)\eta$ is more significant, but the $a_{2}(1320)\pi$ is negligible.
@@
The partial widths obtained in the model generally follow the pattern of $^{3}S_{1}$ models, in which the leading partial wave is much larger than sub-leading ones, as it for all the channels that allow decays into $P$ and $F$ waves.
Figure 2: Plot of $\Gamma(0^{+-}\rightarrow\pi(1300)\pi)$ as a function of the $\pi(1300)$ $\beta$ parameter. The asterisk corresponds to the value of $\Gamma$ at $\beta_{\pi(1300)}=0.167$, which is what we obtain from the spin dependent model in Eq. 11 and report in Table LABEL:tab:IsovectorDecays above.

### v3

@@ Abstract
A model of hybrid meson structure based on the QCD Hamiltonian in Coulomb gauge and a single constituent quasigluon is applied to the computation of hadronic decays of mesons with exotic quantum numbers, $0^{+-}$ and $2^{+-}$. These correspond to hybrid mesons in which the gluon couples to a $q\bar{q}$ pair in an $P$-wave and can therefore be identified as orbital excitations of the exotic $1^{-+}$ state. Interestingly, we find $0^{+-}$ states to be narrow, contrary to what was found by previous calculations. This is primarily due to the suppression of the decay mode $0^{+-}\rightarrow 2^{1}S_{0}+1^{1}S_{0}$, which is unique to our model. The $2^{+-}$ states are found to be narrow, as expected.
@@ I Introduction
Indeed, several experimental efforts such as the GlueX at JLab, PANDA at FAIR and BESIII at BEPC II will be, or are currently, under way and there is hope that future data, combined with continuous progress in lattice field computations, will shed light on these enigmatic states. However, the field is not without progress. The puzzle of the two exotic $1^{-+}$ states labeled $\pi_{1}(1400)$ and $\pi_{1}(1600)$ Workman:2022ynf seems to have mainly been resolved, suggesting that only the heavier of the two is necessary to explain the data JPAC:2018zyd ; Kopf:2020yoa . Recently, the BESIII collaboration BESIII:2022riz has observed an isoscalar resonance with exotic quantum numbers $1^{-+}$ in $J/\psi\to\gamma\eta\eta^{\prime}$. This new state, called $\eta_{1}(1855)$, has a mass of $1855\pm 9{}^{+6}_{-1}$ MeV and a width of $188\pm 18{}^{+3}_{-8}$ MeV, making it an ideal candidate for the isoscalar partner to the isovector $\pi_{1}(1600)$ BESIII:2022riz ; Swanson:2023zlm .
@@
Hybrid mesons models have a long history, dating back to the 1970s, and can be broadly divided into two categories: flux-tube models and constituent gluon models (see Meyer:2015eta for a review). These two camps primarily differ in how they describe the gluonic degrees of freedom. In flux-tube models flux (which include string models string ), glue is represented as a collective string-like excitation, while in constituent models the gluonic degrees of freedom are described by quasiparticles. The latter notion finds confimation in recent lattice studies demonstrating that the low-lying charmonium hybrid spectrum is well-described by a single gluonic quasiparticle with axial quantum numbers $J_{G}^{PC}=1^{+-}$ Dudek:2011bn .
@@
Throughout the years, constituent gluon models have developed from almost dynamics-free models HM to more modern pictures which include many of the features of QCD Swanson:1997wy ; Swanson:1998kx ; Guo:2008yz . More recently, the constituent picture with the QCD Hamiltonian in Coulomb gauge as a starting point has been applied to charmonium hybrid decays Farina:2020slb , the light $H_{1}$ states ($1^{--}$, $(0,1,2)^{-+}$) Farina:2023oqk , and to a study of light hybrid meson mixing Swanson:2023zlm . In this paper, we use this picture to compute the strong decays of the light exotic $0^{+-}$, $2^{+-}$ and $2^{+-^{\prime}}$ hybrid mesons. These correspond to hybrid mesons in which the gluon couples to a $q\bar{q}$ pair in an $P$-wave and can therefore be identified as the orbital excitations of the exotic $1^{-+}$ state previously studied in Farina:2023oqk .
@@
The paper is organized as follows: Section II contains details of the QCD Hamiltonian in Coulomb gauge, which forms the basis of our model, and the construction of the light hybrid meson state from constituent quasiparticles. The model parameters used in the computation and the decay model are also contained in this section, along with the resulting selection rules. Section III presents the results obtained for both isoscalar and isovector hybrids. Summary and conclusions, along with prospects for potential discovery of these states, are found in Section IV.
@@ II.1 QCD Hamiltonian In Coulomb Gauge
The Coulomb kernel is instantaneous and contains all powers of gluonic field operators. These may be organized in a tower of Dyson equations, and truncations can be solved. This approach shows that confinement appears with the aid of a simple variational vacuum ansatzSzczepaniak:2001rg . However, this potential does not necessarily need to be equivalent to the interaction between static color sources, which is typically used in phenomenological calculationsDawid:2024uqz . We therefore choose to simplify by replacing the vacuum matrix element of the Coulomb interaction with:
@@
where $a_{s}$, $b$, and $\mathcal{C}$ are parameters to be determined by fitting to the meson spectrum, as explained below. This choice, besides reproducing the well-known Cornell potential of non-relativistic meson models, finds support in lattice simulations of the $q\bar{q}$ static quark potentials with the gluonic field in different configurations, which show that they all exhibit a long-range linear behavior (see, for example, Ref. Juge:1997nc ).
@@
where $m_{g}$ is the dynamical gluon mass given by $m_{g}\approx 600$ MeV and $b_{g}$ is a parameter with value $b_{g}\approx 6000$ MeV. We remark that despite the appearance of an effective mass, the quasigluon remains transverse.
@@ II.2 Light Hybrids
[…]
The total meson angular momentum $J$ is obtained by first coupling the gluon spin projection to the gluon angular momentum, $\ell_{g}$, to obtain the total gluon angular momentum $j_{g}$. Calculations can be streamlined by working in the gluon helicity basis, which is achieved by means of the relationship
$a^{\dagger,A}_{m_{g}}(\bm{k})=\mathcal{D}^{(1)*}_{m_{g},\lambda}(\hat{k})a^{\dagger,A}_{\lambda}(\bm{k}),$ (8)
where $a^{\dagger}$ is a constituent gluon creation operator, $m_{g}$ is the $z$-component of the gluon angular momentum in the canonical basis and $\lambda$ is the helicity of the gluon. $\mathcal{D}^{(1)*}_{m_{g},\lambda}$ is the Wigner-$\mathcal{D}$ matrix which rotates the operator from the canonical basis to the helicity basis. Doing so and assuming that $\ell_{g}=j_{g}$ produces a Clebsch-Gordan factor of
@@
This represents a transverse electric (TE) gluon and forms the explicit realization of the axial constituent gluon. Setting $\ell_{g}=j_{g}\pm 1$ yields a transverse magnetic (TM) gluon with a Clebsch factor given by $\chi^{(+)}_{\lambda,\mu}=\delta_{\lambda,\mu}/\sqrt{2}$. Combining with quark spins yields the final generic expression for a hybrid ket
@@
The $T^{A}_{ij}$ in the above equation is the color factor, which is obtained by assuming the $q\bar{q}$ to be in a color-octet coupled to the gluon so as to give an overall color singlet, as demanded by color confinement. The explicit form of $\Psi_{j_{g};\ell m_{\ell}}({\bf k,\bf q})$ is given in Eq. 14 below. Since all lowest lying hybrids have $\xi=-1$ and $j_{g}=1$, only the TE mode is of interest for this work. Furthermore, the hybrid state is an eigenstate of parity and charge conjugation with eigenvalues given by
@@ II.3 Model Parameters
The Hamiltonian expectation value of Eq. 13 yields integral expressions in which the angular parts can be done analytically, leaving only radial integrals which can be computed in either momentum or configuration space (for details see Farina:2020slb ). The only matter left to settle is then the choice of basis in which to diagonalize resulting radial system. There are many possible choices and all come with their own advantages and pitfalls. We choose to expand the radial wavefunctions in a finite Gaussian basis of the form $\Psi_{j_{g},\ell}(\vec{k},\vec{q})\propto\Sigma_{n,m}D^{j_{g}*}_{m_{g},\mu}(\hat{k})k\exp(-k^{2}n^{2}/2\beta^{2})\cdot Y_{\ell m_{\ell}}(\hat{q})q^{\ell}\exp(-q^{2}m^{2}/2\alpha^{2})$. This choice is in line with what has been historically done for conventional meson and baryon models, in which authors have used empirical Gaussian expansion methods as an alternative to SHO and hyperspherical methods Richard:1992uk .
The main advantage of this basis is that it yields analytical results for all the integrals in the Hamiltonian expectation value above. Additionally, the expression can be easily Fourier-transformed, allowing the computation in either the configuration or the momentum basis.
Here, however, since the main goal is to compute the decay properties of the mesons, we are only interested in the gross features of meson structure and the resulting hybrid wavefunctions are therefore approximated by a product of two Gaussians (this sets $n=m=1$ in the expansion above) with widths $\alpha$ and $\beta$:
@@
Similarly, ordinary mesons are described by SHO wavefunctions depending on a scale parameter $\beta_{m}$. This parameter is chosen to give the best overlap with the full numerically determined ordinary meson wavefunctions.
@@ II.4 Decay Model
This decay model yields two selection rules that are typical of hybrid decay models, namely: (i) spin zero hybrids do not decay to spin zero mesons, (ii) TE hybrids do not decay to mesons with identical spatial wavefunctions. The former follows from the $^{3}S_{1}$ quantum numbers of the vertex, while the latter follows from the symmetry properties of the decay amplitude (see Ref. Farina:2020slb for details). Additionally, decays to an $S$-wave and $P$-wave meson pair are generally dominant.
@@ III.1 Hybrid Masses
Hybrid masses constitute another source of uncertainty in the model. Because our model only gives spin-averaged hybrid masses, and there are currently no experimental data to constrain them, one must rely on lattice data. However, the results from lattice gauge theory are somewhat scattered. For example, a study of the lightest hybrids supermultiplet Dudek:2011bn computes isovector masses of all the $H_{1}$ hybrids and the exotic $0^{+-}$ and $2^{+-}$ hybrids at different pion masses, ranging from $m_{\pi}\approx 700$ MeV down to $m_{\pi}\approx 400$ MeV. A crude extrapolation down to the physical pion mass gives estimates for the isovector hybrids but not the isoscalars. Alternatively, a more comprehensive study of the full isovector-isoscalar meson spectrum Dudek:2013yja extracted hybrid masses at $m_{\pi}\approx 400$ MeV. The results suggest that there should be a $1^{-+}$ isovector at about $2025$ MeV, and two isoscalar counterparts at higher masses of about $2150$ MeV and $2340$ MeV, respectively. The excited state masses are obtained in the same way, and they show an average mass splitting of about 35 MeV between the isovector and low isoscalar and about 100 MeV between the low and high isoscalar. However, the light isovector $1^{-+}$ mass of about 2025 MeV, is roughly $400$ MeV higher than the accepted $\pi_{1}(1600)$ mass. The approach adopted here, admittedly not overly rigorous, is to shift all of the masses in Ref. Dudek:2013yja down by the same amount (425 MeV) so that the isovector $1^{-+}$ mass matches the $\pi_{1}(1600)$ mass. This keeps the mass splittings the same and gives the masses for the excited states reported in the Table 3.
@@ III.2 Isovectors
$|K_{1L}\rangle$ $=$ $\cos{\theta_{K}}|^{1}P_{1}\rangle+\sin{\theta_{K}}|^{3}P_{1}\rangle$
$|K_{1H}\rangle$ $=$ $-\sin{\theta_{K}}|^{1}P_{1}\rangle+\cos{\theta_{K}}|^{3}P_{1}\rangle.$ (16)
@@
$\Gamma_{tot}$ 22.8 66.9
The results show that the three isovectors are all narrow. While the $2^{+-}$ states were previously reported to be narrow, the $0^{+-}$ comes as a surprise. If we compare our results to those of previous analyses such as the PPS model Page:1998gz (and the IKP model reported therein) or the work of Ref. Chen2025 which uses a very similar approach to ours, the $0^{+-}$ was always reported as being wide, with a total width in the 250-650 MeV range and a dominant contribution of about 200-250 MeV coming from the $\pi(1300)\pi$, which in our work is almost entirely suppressed. In our model, the suppression of the channel is strongly dependent on the value of the $\beta_{m}$ parameter used for the $\pi(1300)$. This dependence is shown in Figure 2, where the total width is plotted as a function of $\beta_{\pi(1300)}$. Oscillatory behavior like the one in the figure are typical of decay channels that include excited states.
Fitting the light ordinary meson spectrum (via Eq. 12) gives $\beta_{\pi}=0.54$ Gev and $\beta_{\pi(1300)}=0.167$ GeV. Indeed, all excited states have relatively small $\beta$ parameters since $\beta$ is inversely proportional to the RMS radius of of meson. Remarkably, the preferred value of $\beta_{\pi(1300)}=0.167$ GeV lies very close to a node in the decay amplitude, which occurs at about $\beta_{\pi(1300)}=0.195$ GeV. The figure also shows that the decay width rapidly increases with increasing $\beta$, reaching its maximum of about $42$ MeV at $\beta_{\pi(1300)}=0.38$ GeV. Thus, even our maximal partial width is smaller than those of Page:1998gz ; Chen2025 by a factor of at least six. It is perhaps useful to point out that, in contrast to our method, authors often employ a single value for the $\beta$ parameters of ground state and excited mesons; as we see here, such an assumption is ill advised.
For all three states, $S+P$ wave decays are preferred. This is the case for the $0^{+-}$, whose main decay channels are the $h_{1}(1170)\pi$ and $b_{1}(1235)\eta$, in that order. The same is also true for the $2^{+-}$ and $2^{+-^{\prime}}$ states. However, these states show some differences. The $2^{+-}$ decays mainly into $a_{1}(1260)\pi$, $a_{2}(1320)\pi$ and $h_{1}(1170)\pi$, in that order, but has a small $b_{1}(1235)\eta$ contribution. In contrast, the main decay channels for the $2^{+-^{\prime}}$ are $h_{1}(1170)\pi$ and $a_{1}(1260)\pi$. In this case, $b_{1}(1235)\eta$ is more significant, but the $a_{2}(1320)\pi$ is negligible.
@@
The partial widths obtained in the model generally follow the pattern of $^{3}S_{1}$ models, in which the leading partial wave is much larger than sub-leading ones, as seen for all the channels that allow decays into $P$ and $F$ waves.
Figure 2: Plot of $\Gamma(0^{+-}\rightarrow\pi(1300)\pi)$ as a function of the $\pi(1300)$ $\beta$ parameter. The asterisk corresponds to the value of $\Gamma$ at $\beta_{\pi(1300)}=0.167$, which is what we obtain from the spin dependent model in Eq. 12 and which is reported in Table LABEL:tab:IsovectorDecays.